OpenStudy (anonymous):

Simplifiy (4x^-2)^-2 with positive exponents

OpenStudy (saifoo.khan):

1 x^4 -- 16

OpenStudy (anonymous):

i believe it would be 4x^4

OpenStudy (anonymous):

nah dood khan is right

OpenStudy (saifoo.khan):

lol @ abstract.

OpenStudy (saifoo.khan):

thnks.

OpenStudy (anonymous):

Can you show your work? Thanks

OpenStudy (saifoo.khan):

Shutto, OK. i can try. Can u join me on twiddla?

OpenStudy (anonymous):

OHHHH lol i wrote down the question wrong. no wonder i couldnt see where i messed up. wrong parenthesis

OpenStudy (saifoo.khan):

Rofl. whts the correct version?

OpenStudy (anonymous):

\[(4x^{-2})^{-2}=4^{-2}x^4\]\[\frac{x^4}{4^2}=\frac{x^4}{16}\]

OpenStudy (anonymous):

I ahve posted step by step explanation: http://www.oojih.com/show/algebra/exponentrationalradical/ scroll down to see your question and the explanation

OpenStudy (anonymous):

Thank you very much